# Deep Learning Foundation and Pattern Models: Challenges in Hydrological Time Series

### v2

@@ 7 Comparison with other Pattern and Foundation Models
Recent advancements in time series analysis have adopted a foundation model approach, drawing inspiration from large language models Mirchandani et al. (2023). Since 2022, there has been a surge in research on foundation models for time series forecasting, primarily utilizing Transformer Vaswani et al. (2023) and MLP Tolstikhin et al. (2021) architectures. Some models incorporate static exogenous variables as input Chen et al. (2023); Das et al. (2023), while others rely solely on time series data Ansari et al. (2024). These models aim to generalize well in data-scarce tasks, integrate multimodal data, and offer explainable decision-making processes Ye et al. (2024). Research in this area focuses on two strategies: pre-training foundation models with time series data and adapting pre-trained language models for time series tasks Ye et al. (2024). Their success stems from effectively capturing multimodal distributions and repeated seasonal patterns Gruver et al. (2024).
@@
This study compares the performance of state-of-the-art foundation models and our LSTM-based pattern model on the hydrology rainfall-runoff task. Pattern models, smaller-scale models trained for specific tasks Jafari et al. (2024), are contrasted with broad-purpose foundation models. All benchmarks use CAMELS-US Addor et al. (2017) data and the Nixtla Neuralforecast framework Olivares et al. (2022). Pattern models include TFT Lim et al. (2021), TCN Bai et al. (2018), DilatedRNN Chang et al. (2017), GNN-Coder Jafari et al. (2024), TimesNet Wu et al. (2022), and TiDE Das et al. (2023). Foundation models analyzed include TSMixer Chen et al. (2023), iTransformer Liu et al. (2023), and PatchTST Nie et al. (2022), pre-trained on datasets like TrafficL Lai et al. (2018), M4 Makridakis et al. (2020), and Weather Wu et al. (2021). Adapted models from LLMs, such as Chronos Ansari et al. (2024), are also evaluated, with Chronos using T5 Raffel et al. (2020) and GPT-2 Radford et al. (2019) backbones.
@@
Results are presented through two experiments: one using min-max and cube root normalization and the other using raw precipitation data.
@@
7.1 Model Performance Comparison of Processed Data Forecasting (Spatial Validation)
Our LSTM model is trained under benchmark configurations in both univariate and multivariate settings, using all 6 time series properties. Missing streamflow values are imputed with the mean of all observed streamflow data for both univariate and multivariate training. We apply min-max and cube root normalization on precipitation and streamflow data. All foundation and pattern models are trained and evaluated on the same training and validation locations, with a sequence length of 21 days and a forecast horizon of 1 day. Static exogenous features are utilized if supported by the model. As shown by the mean squared error results in Table 12, the multivariately trained LSTM model outperforms all other approaches under these benchmark conditions. Note that this study did not study the best hyperparameters that yield the best results for each model.
Model fits shown in Fig. 11, Fig. 11, Fig. 11, and Fig. 11 illustrate LSTM model performance compared to the TSMixer model Chen et al. (2023) that allows static exogenous features along with multivariate time series as input.
Table 12: Performance comparison of various models
Model Type a Input&Output b P c SR d Tmax e Tmin f VP g Q h
LSTM-multivariate P MV&MV 0.00298 0.00259 0.00019 0.00039 0.00015 0.00026
LSTM-univariate P UV&UV 0.00340 0.00321 0.00038 0.00038 0.00024 0.00036
TFT P UV&UV 0.00909 0.00843 0.00077 0.00093 0.00066 0.05686
TSMixer-M4 F MV&MV 0.00721 0.00599 0.00083 0.00051 0.00097 0.00042
TSMixer-TrafficL F MV&MV 0.00738 0.00609 0.00086 0.00053 0.00101 0.00044
TSMixer F MV&MV 0.00866 0.00698 0.00097 0.00073 0.00134 0.00065
TCN P UV&UV 0.01074 0.00889 0.00118 0.00108 0.00147 0.00286
iTransformer-M4 F MV&MV 0.00796 0.00745 0.00125 0.00080 0.00127 0.00041
iTransformer F MV&MV 0.00756 0.00776 0.00127 0.00066 0.00123 0.00054
DilatedRNN P UV&UV 0.00986 0.00896 0.00128 0.00098 0.00141 0.00078
iTransformer-TrafficL F MV&MV 0.00847 0.00790 0.00140 0.00094 0.00139 0.00042
GNNCoder-univariate P UV&UV 0.00710 0.00781 0.00151 0.00119 0.00156 0.00040
PatchTST-Weather F UV&UV 0.00892 0.00847 0.00169 0.00112 0.00164 0.00043
PatchTST-TrafficL F UV&UV 0.00890 0.00847 0.00169 0.00111 0.00165 0.00043
PatchTST-M4 F UV&UV 0.00892 0.00846 0.00169 0.00111 0.00166 0.00043
GNNCoder-multivariate P MV&MV 0.00861 0.00166 0.00171 0.00798 0.00247 0.00040
TimesNet P UV&UV 0.00939 0.00872 0.00176 0.00119 0.00177 0.08661
PatchTST F UV&UV 0.00952 0.00868 0.00184 0.00112 0.00167 0.00055
TiDE P UV&UV 0.01028 0.00870 0.00204 0.00160 0.00268 0.00096
Chronos-T5 F UV&UV 0.03590 0.01527 0.00628 0.00567 0.00724 0.01969
Chronos-GPT2 F UV&UV 0.02804 0.01261 0.00631 0.00663 0.02564 0.01423
Table shows mean squared error results, with the lowest for each predicted target highlighted in red.
Most foundation models prioritize univariate forecasting for broad applicability Ansari et al. (2024). However, a multivariate approach, which jointly trains and forecasts correlated time series properties, can improve performance. Using CAMELS-US data Addor et al. (2017), we found that multivariate training outperformed univariate methods, particularly in tasks like rainfall-runoff modeling where correlations between properties are crucial.
7.2 Model Performance Comparison of Raw Precipitation Forecasting (Temporal Validation)
The evaluation of raw precipitation data reveals notable performance differences among the tested models, as summarized in Table 13 (MSE, MAE, and NNSE).
The MultiFoundationCore achieved the best results (MSE: 26.8345, MAE: 2.4773, NNSE: 0.6312), closely followed by the MultiFoundationPattern (MSE: 26.9578, MAE: 2.5434, NNSE: 0.6305). These advanced architectures effectively model temporal patterns in precipitation, making them robust solutions for hydrological applications.
Among baseline models, LSTM and TFT delivered strong results, with LSTM achieving an NNSE of 0.6305 (MSE: 27.0633, MAE: 2.5089) and TFT slightly behind with an NNSE of 0.6294 (MSE: 27.1544, MAE: 2.3527). Other models, such as TimeLLM (NNSE: 0.6021) and TCN (NNSE: 0.6041), captured temporal dependencies effectively but lagged behind the more advanced architectures.
Foundation models like TSMixer-M4 (MSE: 35.8996) and iTransformer-M4 (MSE: 33.0192) struggled to adapt to raw precipitation data, underscoring the need for domain-specific adjustments. Hybrid models such as DilatedRNN performed competitively (MSE: 27.3794, NNSE: 0.6275), while models like PatchTST-M4 and TiDE showed moderate results, with NNSE values of 0.6096 and 0.6071, respectively, indicating limitations in capturing the complexities of raw precipitation data.
@@ 7.2 Model Performance Comparison of Raw Precipitation Forecasting (Temporal Validation)
Table 13: Performance of time-segmented precipitation forecasting models, where “P” shows pattern models, “F” indicates foundation models, “MV” stands for multivariate inputs/outputs, and “UV” for univariate.
Model Type Input&Output MSE MAE NNSE
TSMixer-M4 F UV&UV 35.8996 3.1405 0.5623
iTransformer-M4 F UV&UV 33.0192 2.8941 0.5827
iTransformer-TrafficL F UV&UV 32.8197 2.7774 0.5842
TSMixer-TrafficL F UV&UV 31.9283 2.8479 0.5909
PatchTST-Weather F UV&UV 31.3607 2.7944 0.5952
TimesNet P MV&UV 30.7948 2.7750 0.5996
TimeLLM F UV&UV 30.4791 2.6698 0.6021
TCN P MV&UV 30.2239 2.6142 0.6041
PatchTST-TrafficL F UV&UV 29.8829 2.7299 0.6068
TiDE P MV&UV 29.8450 2.8262 0.6071
PatchTST-M4 F UV&UV 29.5362 2.7537 0.6096
DilatedRNN P MV&UV 27.3794 2.5771 0.6275
TFT P MV&UV 27.1544 2.3527 0.6294
LSTM P MV&UV 27.0633 2.5089 0.6302
MultiFoundationPattern F MV&UV 26.9578 2.5434 0.6305
MultiFoundationCore F MV&UV 26.8345 2.4773 0.6312
@@
8 MultiFoundationCore Model
The MultiFoundationCore framework enhances time series forecasting in complex scientific domains by integrating multiple pre-trained foundation models and specialized pattern models. Building on the MultiFoundationPattern Jafari et al. (2024), it uses advanced cross-attention mechanisms to synthesize diverse inputs, capturing intricate spatial and temporal dependencies. This scalable approach leverages the strengths of various models to improve prediction accuracy.
@@ 8 MultiFoundationCore Model
9 Conclusion and Future Work
@@
Often, one thinks about deep learning models discovering hidden variables that control observables that one wishes to predict. In this paper, we demonstrate how including environmental factors and exogenous features can improve time series models’ ability to learn hidden variables. We explored CAMELS and Caravan, which together offer large datasets covering many (about 8000) spatial locations and a long time period (around 30 years). Further, there are multiple (6 in CAMELS and 39 in Caravan) observed dynamic streams and up to 209 static (exogenous) parameters for each location. Some of these exogenous and observed data would be natural inputs into a physics model for the catchments. A priori, it is not obvious which exogenous or even endogenous data should be used as they could be implied by the large datasets used in the training and implicitly learned by the training. However, most likely, it is a mixed situation where the combination of observed and exogenous data will give the best results, with the exogenous data declining in importance as the observed dataset grows in size. We clearly show in this paper with the CAMELS and Caravan data that using exogenous data makes significant improvements in the model fits where both static physical quantities and dynamical mathematical functions have value as exogenous information. We emphasize the importance of considering all available data in time series analyses. In future papers, we hope to understand the trade-off better as it varies in the nature and size of exogenous data. We also will present in-depth study of time series pattern and foundation models under varying hyperparameters, using the compilation of hundreds of papers and around 100 models in sci (2023); nix (n.d.); ddz (n.d.); Jin et al. (2023); Liang et al. (2024). We aim to study a range of time series problems to build a taxonomy so that benchmark sets can be built covering the essentially different cases Huang et al. (2019). Through a comprehensive model review, we will also examine univariate versus multivariate time series analysis. Another interesting feature of science time series is that they naturally vary in magnitude by large factors, and this seems not to be very consistent with neural network activation functions at fixed values independent of the stream. We will present a study of new activation functions in the LSTM using a PRELU-like (Parametric Rectified Linear Unit, or PReLU) structure, which naturally deals with magnitude changes in stream values.
@@
Our LSTM model, data preprocessing, and model comparison code is available at https://github.com/JunyangHe/Hydrology. The processed data is available at He (2024).
{acks}
@@
We would like to thank Gregor von Laszewski and Niranda Perera for their contributions and guidance.
{dci}
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
{funding}
The author(s) disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was partially supported by DE-SC0023452: FAIR Surrogate Benchmarks Supporting AI and Simulation Research and the Biocomplexity Institute at the University of Virginia.
{biogs}
Geoffrey Fox received a Ph.D. in Theoretical Physics from Cambridge University, where he was Senior Wrangler. He is now a Professor in the Biocomplexity Institute and Computer Science Department at the University of Virginia. He has an hindex of 90 with over 45,000 citations. He received the High-Performance Parallel and Distributed Computing (HPDC) Achievement Award and the ACM - IEEE CS Ken Kennedy Award for Foundational contributions to parallel computing in 2019. He is a Fellow of APS (Physics) and ACM (Computing) and works on the interdisciplinary interface between computing and applications.
Junyang He is an undergraduate student in computer science at the University of Virginia School of Engineering and Applied Sciences. He is an undergraduate research assistant under the supervision of Professor Geoffrey Fox. His research centers on time series analysis and large-scale spatiotemporal data modeling for scientific applications.
Alireza Jafari is a Ph.D. student in Computer Science at the University of Virginia and a Graduate Research Assistant under the supervision of Professor Geoffrey Fox. He earned his master’s degree in computer science, focusing on graph machine learning from the University of Tehran. Alireza’s research interests include machine learning, time series prediction, foundation models, graph neural networks, and AI applications in science.
Anushka Idamekorala is a Ph.D. student in Computer Science at the University of Virginia and a Graduate Research Assistant under the supervision of Professor Geoffrey Fox. He earned his bachelor’s degree in electronics and telecommunication engineering from the University of Moratuwa, Sri Lanka. Anushka’s research interests include machine learning, time series prediction, foundation models, and AI applications in science.
Ying-Jung Chen received the B.A. in geography from National Taiwan Normal University Taipei city, Taiwan, in 2008 and the M.A, Ph.D. degree in geography from the University of California, Santa Barbara Santa Barbara, CA, in 2018. She is currently pursuing her second master degree in computer science from Georgia Institute of Technology, Atlanta, GA. From 2018 to 2019, She was a joint post-doc Research Associate with the Applied Physics Laboratory and the e-Science Institute from the University of Washington, Seattle, Seattle, WA. Since 2020, She has been a Data and Applied Scientist/Machine Learning Engineer among different industrial sectors, including agriculture, finance,non-profit, and government. Her research interests include time-series analysis/forecasting, AI ethics and safety, remote sensing and geospatial applications. She has served on the program committee forClimate Change AI workshops at NeurIPS 2022-2024, ICLR 2023-2025, and datasafe workshops at AAAI 2025. Finally. she is the reviewer of the journal Water resources management, and Weather, climate, and society.

### v3

@@ VII Comparison with other Pattern and Foundation Models
Recent advancements in time series analysis have adopted a foundation model approach, drawing inspiration from large language models [58]. Since 2022, there has been a surge in research on foundation models for time series forecasting, primarily utilizing Transformer [59] and MLP [60] architectures. Some models incorporate static exogenous variables as input [61, 62], while others rely solely on time series data [63]. These models aim to generalize well in data-scarce tasks, integrate multimodal data, and offer explainable decision-making processes [64]. Research in this area focuses on two strategies: pre-training foundation models with time series data and adapting pre-trained language models for time series tasks [64]. Their success stems from effectively capturing multimodal distributions and repeated seasonal patterns [65].
@@
This study compares the performance of state-of-the-art foundation models and our LSTM-based pattern model on the hydrology rainfall-runoff task. Pattern models, smaller-scale models trained for specific tasks [11], are contrasted with broad-purpose foundation models. All benchmarks use CAMELS-US [27] data and the Nixtla NeuralForecast framework [57]. Pattern models include TFT [66], TCN [67], DilatedRNN [68], GNN-Coder [11], TimesNet [69], and TiDE [62]. Foundation models analyzed include TSMixer [61], iTransformer [70], and PatchTST [71], pre-trained on datasets like TrafficL [72], M4 [73], and Weather [74]. Adapted models from LLMs, such as Chronos [63], are also evaluated, with Chronos using T5 [75] and GPT-2 [76] backbones. For all models, (pattern and foundation) the backbone components are unfrozen during training and fine-tuning. This choice is informed by empirical observations that freezing model backbones leads to a slight drop in performance on CAMELS-US data.
@@
Results are presented through two experiments: one using min-max scaling and cube root transformation and the other using raw precipitation data.
@@
Figure 12: LSTM streamflow fit (temporal validation).
Figure 13: TSMixer streamflow fit (temporal validation).
Figure 14: TimeLLM streamflow fit (temporal validation).
@@
TABLE XIII: Performance of time-segmented precipitation forecasting models, where “P” shows pattern models, “F” indicates foundation models, “MV” stands for multivariate inputs/outputs, and “UV” for univariate.
Model Type Input&Output RMSE NNSE
TSMixer-M4 F UV&UV 5.9916 0.5623
iTransformer-M4 F UV&UV 5.7462 0.5827
iTransformer-TrafficL F UV&UV 5.7288 0.5842
TSMixer-TrafficL F UV&UV 5.6505 0.5909
PatchTST-Weather F UV&UV 5.6001 0.5952
TimesNet P MV&UV 5.5493 0.5996
TimeLLM F UV&UV 5.5208 0.6021
TCN P MV&UV 5.4976 0.6041
PatchTST-TrafficL F UV&UV 5.4665 0.6068
TiDE P MV&UV 5.4631 0.6071
PatchTST-M4 F UV&UV 5.4347 0.6096
DilatedRNN P MV&UV 5.2325 0.6275
TFT P MV&UV 5.2110 0.6294
LSTM P MV&UV 5.2022 0.6302
MultiFoundationPattern F MV&UV 5.1921 0.6305
MultiFoundationCore F MV&UV 5.1802 0.6312
VII-A Model Performance Comparison of Processed Data Forecasting (Spatial Validation)
Our LSTM model is trained under benchmark configurations in both univariate and multivariate settings, using all 6 time series properties. Missing streamflow values are imputed with the mean of all observed streamflow data for both univariate and multivariate training. We apply min-max scaling and cube root transformation on precipitation and streamflow data. All foundation and pattern models are trained and evaluated on the same training and validation locations, with a sequence length of 21 days and a forecast horizon of 1 day. Static exogenous features are utilized if supported by the model. As shown by the root mean squared error results in Table XII, the multivariately trained LSTM model outperforms all other approaches under these benchmark conditions. Note that this study did not study the best hyperparameters that yield the best results for each model.
Model fits shown in Fig. 11, Fig. 11, Fig. 11, and Fig. 11 illustrate LSTM model performance compared to the TSMixer model [61] that allows static exogenous features along with multivariate time series as input.
Most foundation models prioritize univariate forecasting for broad applicability [63]. However, a multivariate approach, which jointly trains and forecasts correlated time series properties, can improve performance. Using CAMELS-US data [27], we found that multivariate training outperformed univariate methods, particularly in tasks like rainfall-runoff modeling where correlations between properties are crucial.
VII-B Model Performance Comparison of Raw Precipitation Forecasting (Temporal Validation)
The evaluation of raw precipitation data reveals notable performance differences among the tested models, as summarized in Table XIII (RMSE and NNSE).
The MultiFoundationCore achieved the best results (RMSE: 5.1802, NNSE: 0.6312), closely followed by the MultiFoundationPattern (RMSE: 5.1921, NNSE: 0.6305). These advanced architectures effectively model temporal patterns in precipitation, making them robust solutions for hydrological applications.
Among baseline models, LSTM and TFT delivered strong results, with LSTM achieving an NNSE of 0.6302 (RMSE: 5.2022) and TFT slightly behind with an NNSE of 0.6294 (RMSE: 5.2110). Other models, such as TimeLLM (NNSE: 0.6021) and TCN (NNSE: 0.6041), captured temporal dependencies effectively but lagged behind the more advanced architectures.
Foundation models like TSMixer-M4 (RMSE: 5.9916) and iTransformer-M4 (RMSE: 5.7462) struggled to adapt to raw precipitation data, underscoring the need for domain-specific adjustments. Hybrid models such as DilatedRNN performed competitively (RMSE: 5.2325), while models like PatchTST-M4 and TiDE showed moderate results, with NNSE values of 0.6096 and 0.6071, respectively, indicating limitations in capturing the complexities of raw precipitation data.
Model fits shown in Fig. 14, Fig. 14, and Fig. 14 illustrate LSTM model performance compared to the TSMixer model [61] and TimeLLM model [77] with temporal validation.
@@ VII-B Model Performance Comparison of Raw Precipitation Forecasting (Temporal Validation)
VIII MultiFoundationCore Model
The MultiFoundationCore framework enhances time series forecasting in complex scientific domains by integrating multiple pre-trained foundation models and specialized pattern models. Building on the MultiFoundationPattern [11], it uses advanced cross-attention mechanisms to synthesize diverse inputs, capturing intricate spatial and temporal dependencies. This scalable approach leverages the strengths of various models to improve prediction accuracy.
@@ VIII MultiFoundationCore Model
IX Conclusion and Future Work
@@
Often, one thinks about deep learning models discovering hidden variables that control observables that one wishes to predict. In this paper, we demonstrate how including environmental factors and exogenous features can improve time series models’ ability to learn hidden variables. We explored CAMELS and Caravan, which together offer large datasets covering many (about 8000) spatial locations and a long time period (around 30 years). Further, there are multiple (6 in CAMELS and 39 in Caravan) observed dynamic streams and up to 209 static (exogenous) parameters for each location. Some of these exogenous and observed data would be natural inputs into a physics model for the catchments. A priori, it is not obvious which exogenous or even endogenous data should be used as they could be implied by the large datasets used in the training and implicitly learned by the training. However, most likely, it is a mixed situation where the combination of observed and exogenous data will give the best results, with the exogenous data declining in importance as the observed dataset grows in size. We clearly show in this paper with the CAMELS and Caravan data that using exogenous data makes significant improvements in the model fits where both static physical quantities and dynamical mathematical functions have value as exogenous information. We emphasize the importance of considering all available data in time series analyses. In future papers, we hope to understand the trade-off better as it varies in the nature and size of exogenous data. We also will present in-depth study of time series pattern and foundation models under varying hyperparameters, using the compilation of hundreds of papers and around 100 models in [6, 7, 8, 9, 10]. We aim to study a range of time series problems to build a taxonomy so that benchmark sets can be built covering the essentially different cases [78]. Through a comprehensive model review, we will also examine univariate versus multivariate time series analysis. Another interesting feature of science time series is that they naturally vary in magnitude by large factors, and this seems not to be very consistent with neural network activation functions at fixed values independent of the stream. We will present a study of new activation functions in the LSTM using a PRELU-like (Parametric Rectified Linear Unit, or PReLU) structure, which naturally deals with magnitude changes in stream values.
@@
Our LSTM model, data preprocessing, and model comparison code is available at https://github.com/JunyangHe/Hydrology. The processed data is available at [79].
Acknowledgment
@@
We would like to thank Gregor von Laszewski and Niranda Perera for their contributions and guidance. We gratefully acknowledge the partial support of DE-SC0023452: FAIR Surrogate Benchmarks Supporting AI and Simulation Research and the Biocomplexity Institute at the University of Virginia.